Let f(x)= √(x-3) ,g(x)= x^2+8

a)Find f(g(x)).
b)Find g(f (x)).

Answer by stanbon(75887)

Let f(x)= √(x-3) ,g(x)= x^2+8
a)Find f(g(x)) = f[x^2+8] = sqrt[x^2+8-3] = sqrt] = sqrt(x^2+5)
------------------------
b)Find g(f(x)) = g(sqrt(x-3)) = ((x-3)+8) = x+5
--------------------
Cheers,
Stan H.
-------------